OpenStudy (anonymous):

Let z= (√3)-i. Express z^5 in both polar form and in the form a+bi.

OpenStudy (anonymous):

you need first \(r=\sqrt{a^2+b^2}\) and then \(\theta\) \[r=\sqrt{\sqrt{3}^2+1^2}=\sqrt{3+1}=2\] \[\theta =\arctan(\frac{-1}{\sqrt{3}})=-\frac{\pi}{6}\] so in polar form it is \[2(\cos(-\frac{\pi}{6})+i\sin(\frac{\pi}{6}))=2e^{-i\frac{\pi}{6}}\]

OpenStudy (anonymous):

\[2(\cos(-\frac{\pi}{6})+i\sin(-\frac{\pi}{6}))=2e^{-i\frac{\pi}{6}}\] typo on that last line

OpenStudy (anonymous):

then to raise to the fifth power, you need only raise \(2^5=32\) and multiply the angles by 5

OpenStudy (anonymous):

you got it from there? clear or no?

OpenStudy (anonymous):

Makes sense:) thanks! i'm a little confused on how to put it into a + bi form though, I'm having trouble with the radical...

OpenStudy (anonymous):

ok it would be much easier of you drew a picture

OpenStudy (anonymous):

but as far as the \(a+bi\) form is conserned, you only need to evaluate sine and cosine at \[-\frac{5\pi}{6}\]

OpenStudy (anonymous):

concerned *

OpenStudy (anonymous):



OpenStudy (anonymous):

you get \[32(\cos(\frac{-5\pi}{6})+i\sin(\frac{-5\pi}{6}))\] \[32(-\frac{\sqrt{3}}{2}+i(-\frac{1}{2})\] \[-16\sqrt{3}-16i\]

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

yw